# Self-Enhancing Video Data Management System for Compositional Events with Large Language Models

### v1

@@ 4.3.2. Data labeling
VOCAL-UDF uses vision-language models (VLMs) to automatically label sampled video frames as positive or negative based on a UDF description. A VLM takes as input an image-text pair and outputs a textual response. However, using a video frame as the image input and the UDF description as a direct query is ineffective, since the concepts we are interested in target specific objects or pairs of objects in the frame, rather than the frame as a whole. Thus, VOCAL-UDF applies the following prompting strategy for attributes and relationships, with the goal of encouraging the VLM to focus on particular objects or interactions between two objects. For attribute concepts, we use the UDF description proposed in Section 4.1 as the text input, and create the image input by sampling an object from the Objects relation and cropping the video frame to include only the object. For relationship concepts, VOCAL-UDF augments the text input with the class names and bounding box coordinates of the relevant objects in the video frame to provide more context to the VLM. It then generates an image patch cropped from the video frame that includes a pair of objects in the same frame from the Objects relation. VOCAL-UDF further augments the image patch by overlaying a red box around the subject and a blue box around the target, thereby providing the VLM with directional information about the relationship. Figure 4 shows an example prompt for labeling the behind relationship.
@@ 4.3.3. Model training
Similar to (Stretcu et al., 2023; Toubal et al., 2024; Daum et al., 2023), VOCAL-UDF leverages a pretrained vision model (e.g., CLIP) as the feature extractor and uses the feature-label pairs to train a multi-layer perceptron (MLP). However, for relationship classification, directly extracting image features from the image patch containing two objects would lead to poor performance, as the feature extractor is not aware of the locations of the objects. Also, many relationships are directional and will not be captured by the feature extractor (e.g., holding is different from holding ). To address these two problems, VOCAL-UDF concatenates image features from three versions of an image patch: the original one containing both objects, one wherein everything except the subject is masked out, and one wherein everything except the target is masked out. For attribute UDFs, VOCAL-UDF simply extracts features from the image patch containing the object. In addition, since we also have the class name of each object from the Objects relation, VOCAL-UDF extracts text features of object classes as input to the MLP as well.
After training an initial model, VOCAL-UDF follows the same approach as (Stretcu et al., 2023; Toubal et al., 2024) to perform several rounds of active learning to select new samples for labeling and retrain the model to further improve its performance. During each active learning iteration, the trained MLP is run over the unlabeled dataset, VOCAL-UDF selects a batch of samples with the highest uncertainty, and these samples are then labeled by the VLM. VOCAL-UDF then retrains the MLP with the updated set of labeled samples.
@@ 4.4. UDF Selection
For a given UDF, VOCAL-UDF can generate both program-based and distilled-model UDFs; when generating program-based UDFs, a set of candidate programs may be implemented. When there are multiple implementation candidates, VOCAL-UDF needs to select the best one. We describe the UDF selection process in this section.
@@ 4.4.1. UDF selection and active learning
When more than one UDF candidate is generated for a proposed UDF, VOCAL-UDF needs to select one. VOCAL-UDF strives to select the UDF that yields the best F1 score for a given set of user labels, as a way of selecting the UDF that best aligns with the user’s intent. To reduce the number of labeled examples needed from the user, VOCAL-UDF uses active learning (Mullapudi et al., 2021; Zhang et al., 2023), which is summarized in Algorithm 1. The algorithm includes several hyperparameters, including a labeling budget , the minimum number of negatives to collect before initiating active learning, and the number of tuples sampled in each iteration. During each iteration, it randomly samples a set of tuples from the database, each containing attribute values corresponding to either one or two objects. If the labeled samples for a class are considered insufficient, VOCAL-UDF selects the sample most likely to belong to that class for labeling (line 1–1). Otherwise, VOCAL-UDF picks the sample with the greatest disagreement among the UDF candidates to help distinguish between them (Zhang et al., 2023) (line 1). More specifically, when selecting a sample that is most likely to belong to a class, VOCAL-UDF priorities the use of VLM annotations over the majority vote of UDF candidates because VLM labels are empirically more reliable than UDF candidates. When selecting a sample with the greatest disagreement, VOCAL-UDF computes a score for a sample of unlabeled data from the database over the set of UDF candidates and then picks the sample with the largest disagreement. The score of each sampled data is computed as the weighted disagreement between the UDF candidates. The weight of each UDF candidate is initialized to (line 1) and set to its performance over the labeled set ( and ) after each iteration (line 1). At the end of each iteration, VOCAL-UDF updates and given the new user labels (line 1).
4.4.2. Dummy UDF
Several factors can contribute to the unsatisfactory performance of generated UDFs, including the complexity of the target concept, the low quality of VLM labels, and an insufficient number of positive training samples for rare events. When no UDF candidate performs well, VOCAL-UDF should not generate a UDF that can hurt the performance and instead should generate a sub-query with the predicate removed. For this, VOCAL-UDF adds a dummy UDF to the list of UDF candidates. The dummy UDF simply returns True and is equivalent to removing the predicate from the query. When all generated UDF candidates perform poorly, VOCAL-UDF can detect this situation and fall back to using the dummy UDF during the UDF selection process.
@@ 4.4.3. UDF generation strategy
Program-based UDFs and distilled-model UDFs offer different trade-offs in terms of performance, interpretability, generation cost and inference throughput. VOCAL-UDF provides the following four strategies based on the user’s preference and the system’s performance requirements: program only generates program-based UDFs, model only generates distilled-model UDFs, llm asks the LLM to decide whether to generate program-based or distilled-model UDFs, and both generates both classes of UDFs. Depending on the user specification, one or more program-based UDFs may be generated for each proposed UDF, and one distilled-model UDF is generated for each proposed UDF. By default, VOCAL-UDF uses the both strategy to maximize query performance. In Section 6.5, we also evaluate the performance of using the llm strategy to automatically choose UDF types.
@@ 5. Implementation
VOCAL-UDF is implemented in Python using the AutoGen framework (Wu et al., 2023). The Query Executor converts queries in DSL notations into SQL and uses a relational engine (DuckDB (Raasveldt and Muehleisen, [n.d.]) in our prototype) to execute them. We apply the same query translation algorithm as in (Zhang et al., 2023) to optimize query execution.
The user comes to VOCAL-UDF with a video dataset, and optionally a set of UDFs. If the set of UDFs includes object detection and tracking models, VOCAL-UDF uses these user-provided models to populate the Objects relation. Otherwise, VOCAL-UDF uses a pre-defined object detection and tracking model to detect and track common objects in the video dataset.
VOCAL-UDF also executes the currently available UDFs ahead of time to populate the Relationships and Attributes relations, so that we have an initial view of the active domains of the oname, rname, and aname columns. The pre-populating happens as soon as VOCAL-UDF receives the videos and UDFs and before the user issues any queries. If the user provides no UDFs, the Relationships and Attributes relations will be empty.
Similarly, when a new UDF is later created, our prototype executes it over the entire dataset and materializes the results in the database. It then converts the UDF into a value-lookup UDF. All UDFs are implemented in Python.
There is preprocessing overhead in the initial setup. The cost of running object detection and tracking models is relatively modest compared to that of LLMs and VLMs, and they can be run on local machines. The cost of executing UDFs depends on their implementation. In this paper, we materialize all results in the video dataset before issuing new user queries. Several works have focused on optimizing the execution of queries with UDFs (Kakkar et al., 2024; Xu et al., 2022), which could be integrated into our system.
@@ 6. Evaluation
Baselines. We compare VOCAL-UDF against VisProg (Gupta and Kembhavi, 2023) and EQUI-VOCAL (Zhang et al., 2023), both of which can only utilize predefined UDFs to answer queries. VisProg uses a different set of modules for different tasks, so we manually write new modules for the system to be able to answer compositional queries, which includes both logical modules (e.g., Eval, Event, Before) and concept modules (e.g., Red, Holding). VisProg also requires in-context examples, which we create separately from the evaluation queries. EQUI-VOCAL does not need such examples. EQUI-VOCAL also does not handle NL queries, so we add a query interface that converts NL to the EQUI-VOCAL DSL using an LLM, which is the same as VOCAL-UDF.
@@
Metrics. We evaluate query answering performance using F1 scores. We propose and generate UDFs using training data and report query F1 scores over the test data. We evaluate each dataset using 30 queries, and each query is run three times.
Evaluation setup. We conduct all experiments on a compute cluster. For each experiment, we request one node with eight Intel Xeon Gold 6230R CPUs at 2.10GHz, 200GB of RAM, and one NVIDIA A40 GPU. We use the GPT-4 Turbo model (gpt-4-turbo-2024-04-09) as the LLM and VLM across all systems. We configure VOCAL-UDF as follows. For the CLEVRER dataset, we use a labeling budget of 20 for each UDF during selection. In program-based UDF generation, 10 candidate programs are generated, and numeric parameters and frame pixels are allowed as function inputs. Each parameter is then instantiated with the default value as well as five random values during the UDF selection stage. When generating distilled-model UDFs, we ask the VLM to annotate 100 sampled frames per UDF. For the CityFlow-NL and Charades datasets, we set the labeling budget of each UDF selection to 50 and the number of VLM-annotated frames to 500 per UDF, since the datasets are more challenging. Frame pixels are disallowed during program-based UDF generation due to the minimal performance improvement it offers compared to the substantial increase in execution time. We assume objects are already detected and tracked, so we can focus our evaluation on relationships and attributes.
@@
Datasets: CLEVRER: The CLEVRER (Yi et al., 2020) dataset consists of 10,000, 5-second synthetic videos of moving objects. To determine ground-truth information, following (Yi et al., 2020), we use a Mask R-CNN (He et al., 2017) to locate objects and predict their colors, shapes, and materials, since the original dataset does not provide bounding box information for objects. We consider eight colors, two material types, and three shapes. Further, we write rule-based functions to extract spatial relationships (Near, Far, LeftOf, RightOf, FrontOf, Behind) and attributes (Left, Right, Top, Bottom). For our experiments, we use the same Mask R-CNN and rule-based functions as UDFs, for a total of six relationships and 17 attributes. Among them, we select two relationships and nine attributes as the base UDFs (the user-provided UDFs available to all systems during query evaluation). We use templates to automatically generate 30 target queries with seven predicates (three of which are from the supplemental UDFs, which are the UDFs that will need to be generated, and the others are from the base UDFs), up to three variables (i.e., three distinct objects), up to three region graphs, and duration constraints with three possible values. We ensure that there are at least 5% positive examples in the dataset for each target query (we ensure the same for all datasets). We then rewrite the DSL queries as NL queries using GPT-4 and manually verify the correctness (translating NL to DSL is not a contribution of this paper). To generate ground truth labels, we run each target DSL query on the dataset.
@@
CityFlow-NL: The CityFlow-NL (Feng et al., 2021) dataset contains traffic videos captured from multiple cameras and NL descriptions for vehicle tracks. Following (Le et al., 2023), we extract colors and types of vehicles from the NL descriptions. Since annotations are supplied only for sampled vehicle tracks, we only consider annotated tracks in our evaluation. We extract six spatial relationships (Above, Beneath, LeftOf, RightOf, FrontOf, Behind) using rule-based predictors. We create 1473 non-overlapping 50-frame video segments from the original dataset. To create UDFs for attributes, we train binary image classifiers. We use half of the videos as training data and the rest as test data. For relationships, we use the same rule-based functions as UDFs. Then, we select three relationships and four attributes as base UDFs. We automatically generate 30 target queries with up to four predicates (two of which are from the supplemental UDFs), up to three variables, up to three region graphs, and duration constraints with three possible values.
@@
Charades: The Charades (Ji et al., 2020) dataset contains 30-second videos of daily indoor activities. Action Genome (Ji et al., 2020) provides scene graph annotations for Charades on selected frames. In the evaluation, we focus solely on these annotated frames, which preserves the relative temporal order of the frames but means we cannot determine if the frames are contiguous. Consequently, duration constraints are not considered in the queries. The scene graph annotations encompass humans, objects, and human-object relationships, but do not cover attributes or object-object relationships. Thus, our evaluation queries focus on humans’ interactions with objects in the scene. At query time, only human-object pairs are sampled to collect more positive examples. Since Action Genome is annotated by human labelers, the annotations are incomplete, which is a common issue in real-world datasets (Chen et al., 2024). We therefore augment the dataset by automatically generating dense spatial relationship annotations using rule-based predictors to replace the existing annotations for spatial relationships. In total, we have 20 relationships (five spatial, 15 semantic) and 35 object classes. Among them, we select two spatial relationships and seven semantic relationships as the base UDFs. When generating queries, we need to restrict predicates to human-object relationships. We automatically generate 30 target queries with up to four predicates (one is always object(oi,‘person’), two are from the supplemental UDFs), up to three variables, and up to two region graphs, but without any duration constraints.

### v2

@@ 4.3.2. Data labeling
VOCAL-UDF uses VLMs to automatically label sampled video frames as positive or negative based on a UDF description. A VLM takes as input an image-text pair and outputs a textual response. However, using a video frame and the UDF description as a direct query is ineffective, since the concepts we are interested in target specific objects or pairs of objects, rather than the frame as a whole. Thus, VOCAL-UDF applies the following prompting strategy, with the goal of encouraging the VLM to focus on particular objects or interactions between two objects. For attribute concepts, we use the UDF description proposed in Section 4.1 as the text input, and create the image input by sampling an object from the Objects relation and cropping the video frame to include only the object. For relationship concepts, VOCAL-UDF augments the text input with the class names and bounding box coordinates of the relevant objects in the video frame to provide more context to the VLM. It then generates an image patch cropped from the video frame that includes a pair of objects in the same frame from the Objects relation. VOCAL-UDF further augments the image patch by overlaying a red box around the subject and a blue box around the target, thereby providing the VLM with directional information about the relationship. Figure 4 shows an example prompt for labeling the behind relationship.
@@ 4.3.3. Model training
Similar to (Stretcu et al., 2023; Toubal et al., 2024; Daum et al., 2023), VOCAL-UDF leverages a pretrained vision model (e.g., CLIP) as the feature extractor and uses the feature-label pairs to train a multi-layer perceptron (MLP). However, for relationship classification, directly extracting features from the image patch containing two objects would perform poorly, as the feature extractor is not aware of the object locations and cannot capture directional relationships (e.g., holding vs. holding ). To address this, VOCAL-UDF concatenates features from three versions of an image patch: the original one containing both objects, one where everything except the subject is masked, and one where everything except the target is masked. For attribute UDFs, VOCAL-UDF simply extracts features from the image patch containing the object. In addition, VOCAL-UDF incorporates text features of object class names, derived from the Objects relation, to enhance the MLP’s performance. After training an initial model, VOCAL-UDF uses an active learning approach similar to (Stretcu et al., 2023; Toubal et al., 2024) to iteratively improve the model. During each iteration, the trained MLP is run over the unlabeled dataset, VOCAL-UDF selects a batch of samples with the highest uncertainty for labeling by the VLM. VOCAL-UDF then retrains the MLP with the updated labeled dataset.
VOCAL-UDF uses VLMs to automatically label data and the cost is bounded by the training data size, which is at most 500 frames in our evaluation. This is only 0.04% of the 1.28M frames in the entire database for CLEVRER, 0.8% of the 65K frames for CityFlow-NL, and 0.2% of the 289K frames for Charades (see Section 6), which correspond to extremely low sampling rates. This cost is thus far lower than directly applying an LLM to the entire video dataset, even with downsampling. Although VOCAL-UDF’s distilled model process may occasionally deliver unsatisfactory performance, its UDF Selector component, described next, mitigates this by selecting the best candidate UDFs and filtering low-performing models.
@@ 4.4. UDF Selection
For a given UDF, VOCAL-UDF can generate both a distilled-model and a set of program-based implementations. When there are multiple implementation candidates, VOCAL-UDF needs to select the best one. VOCAL-UDF does not have any initial data to validate semantic correctness, only the user NL query. If VOCAL-UDF were to collect labeled data for validation before query execution, the user would need to manually go through the video dataset to find positive and negative examples. Instead, VOCAL-UDF utilizes active learning at query time to strategically request binary labels on carefully selected samples that are most likely to resolve disagreements among UDF candidates. Additionally, VOCAL-UDF leverages LLMs to generate the UDF Python code on behalf of the user (Section 4.2) to reduce user effort. Manually writing code snippets for a UDF would require knowledge of Python and could be a tedious trial-and-error process for the user, especially when a UDF requires hyperparameters, operates on pixel values, or leverages existing UDFs as building blocks. We now describe the UDF selection process in detail.
@@ 4.4.1. UDF selection and active learning
To best align with the user’s intent, VOCAL-UDF strives to select the candidate UDF that yields the best F1 score for a given set of user labels. As summarized in Algorithm 1, VOCAL-UDF uses active learning (Mullapudi et al., 2021; Zhang et al., 2023b; Karimi et al., 2021) to reduce the number of labeled examples needed from the user. This algorithm’s hyperparameters include: a labeling budget , thresholds and giving the number of positive and negative samples needed to initiate active learning, and the number of tuples sampled in each iteration. Section 6 details the values used in our experiments.
During each iteration, the algorithm randomly samples tuples from the database (Line 1). It operates in two phases: bootstrapping and active learning. The bootstrapping phase collects at least positive and negative samples. While the number of labeled positives () or negatives () remains below its respective threshold, VOCAL-UDF selects samples most likely to be positive or negative (Lines 1 to 1). Once enough initial labeled data has been collected, VOCAL-UDF transitions to active learning, adopting the margin and positive strategy from (Mullapudi et al., 2021). Specifically, if , it asks the user to label samples most likely to be positive; otherwise, it picks the sample with the greatest disagreement among UDF candidates to differentiate them (Zhang et al., 2023b) (Lines 1 to 1).
In and , the algorithm priorities VLM annotations obtained in Section 4.3.2 over the UDF candidates’ majority vote, as VLM labels are empirically more reliable. In , it computes a disagreement score for each sample in based on UDF candidates , selecting the one with the highest disagreement (Karimi et al., 2021; Zhang et al., 2023b). Each sample’s disagreement score is the weighted disagreement among UDF candidates, with candidate weights initialized to (Line 1) and updated to its performance (F1 score in our prototype) over and after each iteration (Line 1). At the end of each iteration, , and are updated with the new user labels (Line 1).
4.4.2. Dummy UDFs
Several factors can contribute to the unsatisfactory performance of generated UDFs, including the complexity of the target concept, the low quality of VLM labels, and an insufficient number of positive training samples for rare events. When no UDF candidate performs well, VOCAL-UDF should not apply a UDF that can hurt the performance. To prevent this, VOCAL-UDF appends a dummy UDF to its list of UDF candidates. A dummy UDF is a constant-valued function that produces True and is equivalent to omitting the predicate. When all other candidates perform poorly, VOCAL-UDF opts for the dummy UDF during its UDF selection process.
@@ 4.4.3. UDF generation strategy
Program-based UDFs and distilled-model UDFs offer different trade-offs in terms of performance, interpretability, generation cost, and inference throughput. VOCAL-UDF provides the following four strategies based on the user’s preference and the system’s performance requirements: program only generates program-based UDFs, model only generates distilled-model UDFs, llm asks the LLM to decide whether to generate program-based or distilled-model UDFs, and both generates both classes of UDFs. LABEL:fig:deciding_udf_type_prompt shows an example prompt of using the llm UDF generation strategy. Depending on the user specification, one or more program-based UDFs may be generated for each proposed UDF, and one distilled-model UDF is generated for each proposed UDF. By default, VOCAL-UDF uses the both strategy to maximize query performance. In Section 6.4, we also evaluate the performance of using the llm strategy to automatically choose UDF types.
4.5. Comparing with direct LLM methods
As an alternative to VOCAL-UDF, a multimodal LLM can directly answer user queries. However, as shown in prior work (Ma et al., 2023; Wang et al., 2024) and as we confirm in Section 6.1.2, LLMs still struggle with compositional queries, and directly providing the LLM with the video frames and the user query in a single prompt (the “vanilla LLM” approach in Section 6.1.2) leads to low F1 scores. Furthermore, the cost scales with the dataset size, making it expensive for large video collections today, as we discussed in Section 4.3 and evaluate in Table 3.
A higher performance LLM baseline is to parse the query into our DSL, then use a combination of existing cheap models for common concepts and an LLM for application-specific concepts (i.e., VOCAL-UDF without UDF generation, or “LLM for concepts” in Table 3). While it may yield good F1 scores, this approach is only practical for short videos and a small number of queries. For larger workloads, the cost and latency of this method are even greater than “vanilla LLM” due to repeated LLM calls to identify concepts for each object or object pair. As Table 3 shows, executing just one query with three missing UDFs over the CLEVRER dataset using GPT-4o costs about $1,504 USD. Reducing costs through sampling is also non-trivial because low sampling rates can degrade query performance, especially for compositional events involving temporal dimensions. As we show in Figure 6(b), even a small downsampling factor of 8 can greatly reduce the upper bound of achievable F1 scores from 1.0 to 0.51.
VOCAL-UDF overcomes these issues by using the multimodal LLM to learn each new application-specific concept once, over a fixed-size sample of the data, and generate cheaper, domain-specific models (either program-based UDFs or distilled-model UDFs) that can then be applied to the remainder of the videos and to all future queries, avoiding repeated reliance on the LLM and reducing overall costs.
@@ 5. Implementation
VOCAL-UDF is implemented in Python using the AutoGen framework (Wu et al., 2023). The Query Executor converts queries in DSL notation into SQL and uses a relational engine (DuckDB (Raasveldt and Muehleisen, [n. d.]) in our prototype) to execute them. We apply the same query translation algorithm as in (Zhang et al., 2023b) to optimize query execution. To reduce system latency, we parallelize API calls to the LLM.
VOCAL-UDF requires a preprocessing stage per dataset before queries can be issued. The user provides a video dataset and optionally a set of UDFs. If these UDFs include object detection and tracking models, VOCAL-UDF uses them to populate the Objects relation. Otherwise, VOCAL-UDF uses a predefined object detection and tracking model to identify common objects. VOCAL-UDF pre-extracts features into Parquet files for use by distilled models at query time (Section 4.3.3), and we utilize NVIDIA DALI (nvi, 2024) to accelerate this process. VOCAL-UDF also executes the initial UDFs ahead of time to populate the Relationships and Attributes relations and compute the active domains for oname, rname, and aname. These operations happen as soon as VOCAL-UDF receives the videos and UDFs and before the user issues any queries. If no initial UDFs are supplied, the Relationships and Attributes relations will be empty. Similarly, when a new UDF is later created, our prototype runs it over the entire dataset and materializes the results in the database. It then converts the UDF into a value-lookup UDF so that subsequent queries can reuse the materialized results. All UDFs are implemented in Python.
There is preprocessing overhead in the initial setup. The cost of running object detection, object tracking, and feature extraction is relatively modest compared to that of LLMs and VLMs, and they can be run locally. The cost of executing a UDF depends on its implementation. In this paper, we materialize all results in the video dataset before issuing new queries. Several works have focused on optimizing the execution of queries with UDFs (Kakkar et al., 2024; Xu et al., 2022), which could be integrated into our system.
@@ 6. Evaluation
Baselines. We compare VOCAL-UDF against VisProg (Gupta and Kembhavi, 2023) and EQUI-VOCAL (Zhang et al., 2023b), both of which are limited to predefined UDFs when answering queries. VisProg uses a different set of modules for different tasks, so we manually write new modules for the system to be able to answer compositional queries. These include both logical (e.g., Eval, Event, Before) and conceptual (e.g., Red, Holding) modules. VisProg also requires in-context examples, which we create separately from the evaluation queries. While EQUI-VOCAL does not need such examples, it does not accept NL queries as input. To address this, we adapt VOCAL-UDF’s approach of converting NL into the EQUI-VOCAL DSL using an LLM. We further compare VOCAL-UDF with two LLM baselines as described in Section 4.5.
@@
Metrics. We evaluate query answering performance using F1 scores, precision, and recall. We propose and generate UDFs using training data and report query performance over the test set. We evaluate each dataset using 30 queries, and each query is run three times. We additionally evaluate the system efficiency in terms of latency and cost. For latency, we report the wall-clock time as perceived by the user from the moment they submit a query to the moment when VOCAL-UDF returns the query results. For cost, we measure both the LLM API invocation cost using the OpenAI pricing model (ope, 2024) and the resource cost using the AWS pricing model (aws, 2024b, a).
Datasets: We evaluate VOCAL-UDF on three datasets: CLEVRER (Yi et al., 2020), CityFlow-NL (Feng et al., 2021) and Charades (Ji et al., 2020). CLEVRER is a benchmark that facilitates testing queries with varying complexities. CityFlow-NL and Charades represent real-world applications—traffic monitoring and human activities—typical for video data management systems. UDFs in CityFlow-NL include fine-grained vehicle types, while Charades comes with complex human interactions. The evaluation queries are based on concepts from these datasets, which are labeled by the original annotators and reflect events from those domains.
@@
CLEVRER: The CLEVRER (Yi et al., 2020) dataset consists of 10,000, 5-second synthetic videos of moving objects. To determine ground-truth information, following (Yi et al., 2020), we use a Mask R-CNN (He et al., 2017) to locate objects and predict their colors, shapes, and materials. We write rule-based functions to extract spatial relationships and attributes. For our experiments, we use the same Mask R-CNN and rule-based functions as UDFs, for a total of six relationships and 17 attributes. Among them, we select two relationships and nine attributes as the base UDFs, which are available to all systems during query evaluation. We use templates to automatically generate 30 target queries with seven predicates, up to three variables (i.e., three distinct objects), up to three region graphs, and duration constraints with three possible values. We ensure there are at least 5% positive examples in the dataset for each target query (we ensure the same for all datasets). We then rewrite the DSL queries as NL queries using GPT-4 and manually verify the correctness.
@@
CityFlow-NL: The CityFlow-NL (Feng et al., 2021) dataset contains traffic videos captured from multiple cameras and NL descriptions for vehicle tracks. Following (Le et al., 2023), we extract vehicle colors and types from the NL descriptions. Since only sampled vehicle tracks are annotated, we only consider them in our evaluation. We extract six spatial relationships using rule-based predictors and create 1,473 non-overlapping 50-frame video segments from the original dataset. To create UDFs for attributes, we train binary image classifiers. For relationships, we use the same rule-based functions as UDFs. Then, we select three relationships and four attributes as base UDFs. We automatically generate 30 queries with up to four predicates, up to three variables, up to three region graphs, and duration constraints with three possible values. An example query is ”a vehicle o0 with type sedan is to the right of another vehicle o1 for at least 15 frames. Then, o0 is to the left of o1 for at least 15 frames”, which could be used to find sedans changing their positions with other vehicles.
@@
Charades: The Charades (Ji et al., 2020) dataset contains 30-second indoor activity videos. We focus exclusively on frames with scene graph annotations provided by Action Genome (Ji et al., 2020), which only provide human-object relationships but do not cover attributes or object-object relationships. Thus, our evaluation queries focus on human interactions with objects. At query time, only human-object pairs are sampled to collect more positive examples. Incomplete annotations is a common issue in real-world datasets (Chen et al., 2024d) like Charades. To address this, we augment the dataset by automatically generating dense spatial relationship annotations using rule-based predictors to replace the existing ones. For semantic relationships, we train binary image classifiers. In total, we have 20 relationships (five spatial, 15 semantic) and 35 object classes. Among them, we select two spatial and seven semantic relationships as the base UDFs. We automatically generate 30 target queries with up to four predicates (one is always object(oi,‘person’)), up to three variables, and up to two region graphs, but without duration constraints. An example query is ”a person is eating something, and then the same person is holding another object while being in front of it”, which could retrieve videos including a person taking a bite from an apple, and then picking up a cookbook in front of them to read a recipe.
For each dataset, we use half of the videos as training data and the rest as test data. To generate ground truth labels, we run each target DSL query on the dataset.
Evaluation setup. We conduct all experiments on a compute cluster. For each experiment, we request one node with eight Intel Xeon Gold 6230R CPUs at 2.10GHz, 200GB of RAM, and one NVIDIA A40 GPU. We use the GPT-4o model (gpt-4o-2024-08-06) as the LLM and VLM across all systems. We configure VOCAL-UDF as follows. For the CLEVRER dataset, we use a labeling budget of 20 for each UDF during selection. We generate 10 candidate programs, allowing numeric parameters and frame pixels as inputs. During UDF selection, we bind each numeric parameter to its default value as well as five random values. When generating distilled-model UDFs, we ask the VLM to annotate 100 sampled frames per UDF. Given their increased complexity, for the CityFlow-NL and Charades datasets, we set the labeling budget per UDF to 50 and the number of VLM-annotated frames per UDF to 500. 100 frames are annotated for the initial training, and then active learning is performed to select 100 additional frames in each round. Frame pixels are disallowed during program-based UDF generation due to their limited impact on performance. To focus our evaluation on relationships and attributes, we assume objects have already been detected and tracked.